OpenStudy (khantahmina):

xy-3x^2=y implict difference

OpenStudy (khantahmina):

@triciaal @steve816

OpenStudy (sweetburger):

\[\frac{ d }{ dx }[xy]+\frac{ d }{ dx }[-3x^2]=\frac{ d }{ dx }[y]\]

OpenStudy (sweetburger):

The hardest part is using the product rule on the first term. Do you know how to?

OpenStudy (khantahmina):

no :(

OpenStudy (khantahmina):

this is something i got \[x (dy/dx) +y-6x=1 dy/dx\]

OpenStudy (sweetburger):

yes, \[x \frac{ dy }{ dx }+y-6x=\frac{ dy }{ dx }\] is correct so far. You need to isolate \(\frac{ dy }{ dx }\)

OpenStudy (khantahmina):

-y+6x ?

OpenStudy (khantahmina):

how would i isloate it

OpenStudy (phi):

it's easier to type y' for dy/dx move all the y' terms to one side, and everything else to the other side

OpenStudy (sweetburger):

Agreed, I thought showing the operator might be helpful.

OpenStudy (khantahmina):

i yet dont know what to do

OpenStudy (khantahmina):

ok i got it \[dy/dx= (6x-y)/(x-1)\]

OpenStudy (khantahmina):

thanks